Q: Find the local and absolute extreme values of the function on the given interval. Also

specify the intervals where function is increasing or decreasing

𝑓(π‘₯) = π‘₯2e-x





1
Expert's answer
2022-01-10T16:33:01-0500
"f(x)=x^2e^{-x}"


Domain: "(-\infin, \infin)"

Find the fist derivative with respect to "x"

"f'(x)=(x^2e^{-x})'=2xe^{-x}-x^2e^{-x}"

Find the critical number(s)


"f'(x)=0=>2xe^{-x}-x^2e^{-x}=0"

"xe^{-x}(2-x)=0"

"x_1=0, x_2=2"

Critical numbers: "0, 2."

Find the second derivative with respect to "x"

"f''(x)=(2xe^{-x}-x^2e^{-x})'"




"=2e^{-x}-2xe^{-x}-2xe^{-x}+x^2e^{-x}"

"=e^{-x}(2-4x+x^2)"

"f''(0)=e^{-0}(2-4(0)+(0)^2)=2>0"

"f''(0)=e^{-2}(2-4(2)+(2)^2)=-2e^{-2}<0"

i) If "x\in [-1, 3]"



"f(-1)=(-1)^2e^{-(-1)}=e"

"f(3)=(3)^2e^{-3}=9e^{-3}"

"f(0)=(0)^2e^{-0}=0"

"f(2)=(2)^2e^{-2}=4e^{-2}"

The function "f(x)" has a local maximum with value of "4e^{-2}" at "x=2."

The function "f(x)" has a local minimum with value of at "x=0."

The function "f(x)" has the absolute maximum on "[-1, 3]" with value of "e" at "x=-1."

The function "f(x)" has the absolute minimum on "[-1, 3]" with value of at "x=0."


ii)

If "x<0, f'(x)<0, f(x)" decreases.

If "0<x<2, f'(x)>0, f(x)" increases.

If "x>2, f'(x)<0, f(x)" decreases.

The function "f(x)" increases on "(0, 2)."

The function "f(x)" decreases on "(-\infin, 0)\cup(2, \infin)."